OpenStudy (anonymous):

Solve the following equation, answer as a reduced, mixed number . Then place the correct number in the box provided. 3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1)

OpenStudy (texaschic101):

3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1) -- distribute through the parenthesis 3x - 6 + 8x - 24 = 6x - 24 + 16x + 8 -- combine like terms 11x - 30 = 22x - 16 -- subtract 22x from both sides 11x - 22x - 30 = -16 -- add 30 to both sides 11x - 22x = 30 - 16 -- combine like terms -11x = 14 -- divide by -11 x = -14/11 x = - 1 3/11